If 3 sides of a triangle are 3x+4y, 4x+3y and 5x+5y (x,y>0) then this traingle is:
 
A EQuilateral
B Right Angled
C  Obtuse angled
D Isoceles

Grade:11

1 Answers

jagdish singh singh
8 years ago
\hspace{-0.6cm}$ Angle opposite to largest side of a triangle are largest. So Using\\\\\\ $\cos C=\frac{a^2+b^2-c^2}{2ab}\;,$ Where $a=3x+4y\;,b=4x+3y\;,c=5x+5y$\\\\\\and $x,y>0$ and Here $C$ is a largest angle.\\\\\\So $\cos C= \frac{-2xy}{(3x+4y)\cdot(4x+3y)}<0\Rightarrow C=\bf{obtuse\; angle}\;,$ $x,y>0$